The cost of building an office complex, x floors high, in a prime location in Accra is made up of three components:

 (a) GH¢10 million for the land

(b) GH¢1/4 million per floor

(c) Specialized costs of GH¢10000x per floor.

 How many floors should the office complex contain if the average cost per floor is to be minimized?


1
Expert's answer
2021-05-19T10:49:34-0400

4. Average cost per floor ATC is minimized if

"ATC = MC."


"\frac{1}{4}=0.25"


"ATC = \frac{TC}{x }=\frac{ (10 + 0.25x)}{x }=\frac {10}{x }+0.25"


"MC = TC'(x) = 0.25"


"\frac{10}{x} + 0.25 = 0.25"


"\frac{10}{x} = 0."


There is no root, so, there is no point, at which ATC is minimized.